OpenStudy (anonymous):

Simplify the expression. (5 + 2)[6 – (3 + 2)]

HanAkoSolo (jamierox4ev3r):

Hi @jordanjamesbay , first add the values in the parentheses farthest to the right.

OpenStudy (anonymous):

Parenthesis Exponent Multiplication Division Addition Subtraction

OpenStudy (anonymous):

7?

HanAkoSolo (jamierox4ev3r):

then slowly work your way towards the right (5+2)[6-(3+2)] (5+2)[6-(5)] (5+2) x 1

OpenStudy (anonymous):

Yes

HanAkoSolo (jamierox4ev3r):

Yes :)

OpenStudy (anonymous):

lol that was easy

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

(5+2) = 7 (3+2) = 5 7(6-5) (6-5) = 1 7* 1 = 7